Kentucky Section of MAA Annual Meeting Student Registration Form

Bellarmine University, Louisville April 4 - 5, 2003

Name: School: E-mail:

I am:  Fr  So  Jr  Sr  Grad

Check all that apply:

I would like to come to dinner on Friday evening. I could use some help finding a place to sleep. I would like to give a talk on Friday afternoon. I would like to give a talk on Saturday morning. I would like to attend the MAA business luncheon on Saturday afternoon.

Please enclose $ 7 for lunch

payable to the KY section of the MAA.

If you would like to give a talk at the meeting: submit the Presentation Abstract Form to Bill Fenton by March 3, 2003.

The information on this form must be received by March 20th or at on-site registration. Submit to: Dora Ahmadi, Dept of Math, Morehead State University,

Morehead, KY 40351 Phone: (606) 783 - 2919

E-mail submissions are encouraged! E-mail: d.ahmadi@morehead-st.edu